Math. 312 Schedule

(Except for test days, this is only approximate.)

Block 1:

Topic  Days Haberman sections  Fulling notes pages
Introductory examples*  5  1.1–1.4, 2.1–2.4  1–14
Fourier series  4  3.1–3.3, 3.6  15–28
Linearity and homogeneity  1  2.2  29–35
Catchup or review  1
Test A  Friday, Sept. 22

* wave equation in lecture, heat equation in textbook Block 2:

Topic  Days Haberman sections  Fulling notes pages
Rectangle problems  1  2.5.1, 7.1–7.4  36–42
Fourier transforms  4  10.1–10.6  43–54
Green functions  5  9.1–9.3, (11.3)*  55–71
Catchup or review  1
Test B  Friday, Oct. 20

* Parts of Sec. 11.3 assume that you’ve studied all of Chapters 8 and 9. Block 3:

Topic  Days Haberman sections  Fulling notes pages
Sturm–Liouville problems  4  5.1–5.5, 5.8–5.10, 7.5 72–81
Polar coords., Bessel fns.  5  1.5, 2.5.2, 7.7–7.9  82–101
Catchup or review  2
Test A  Friday, Nov. 17

Block 4:

Topic  Days Haberman sections  Fulling notes pages
Spherical harmonics  3  7.10  102–109
Classification  2  2.5.4, 6.1  110–113
Catchup or review for final  1

Advanced reading:

Topic Haberman sections Fulling notes pages
More on the wave equation  Ch. 4, 11.2, Ch. 12
Green fns. for nonhom. problems  Ch. 8, 9.5
Convergence thms. for Fourier series  3.4–3.5  App. B